"A carpenter sells a center table for Rs 322, getting (1/6) of its cost price. Find his gain percent."
Answer by Theo(13342)

he sells the center table for 322.
his profit is (1/6) of the cost.
the general equation is:
p = r - c
this means profit equals revenue minus cost.
his revenue is 322.
his profit is (1/6) times the cost.
the equation becomes:
(1/6)c = 322 - c
add c to both sides of this equation to get:
c + (1/6)c = 322
combine like terms to get:
(7/6)*c = 322
multiply both sides of this equation by (6/7) to get:
c = 322 * (6/7) which makes c equal to 276
the cost is 276
the selling price is 322
the profit is 322 - 276 = 46
ratio of profit to cost is 46 / 276 which can be simplified to 1/6.